OpenStudy (ledah):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 3x + 2.

OpenStudy (anonymous):

3(-2)+2 = -4 3(0)+2= 2 3(2)+2 = 8 Range: \( (-4,2,8) \)

OpenStudy (ledah):

I got the opposite of that 8,2-4 :/

OpenStudy (ledah):

could you answer another question for me please?

OpenStudy (anonymous):

That's the same thing, order doesn't matter in sets.

OpenStudy (anonymous):

In more general notation the answer is \( \{ -4,2,8 \} \)

OpenStudy (ledah):

Put the equation in slope-intercept form and identify the slope of the line. 3x + 9y = 12 (Nobody else answered this when i had it posted)

OpenStudy (ledah):

I dont quite know how to work this out

OpenStudy (anonymous):

\(3x + 9y = 12\) This is a equation of straight line. \( 3x + 9y = 12 \implies y = -\frac 13x+\frac 4 3 \) This is slope intercept form, where slope is \(-\frac 13\)

OpenStudy (ledah):

Ah. alright thanks @FoolForMath

OpenStudy (anonymous):

Glad to help :)